A 50 year old patient is hospitalized for knee surgery. Post surgical, the patient rates her...

  1. A 50 year old patient is hospitalized for knee surgery. Post surgical, the patient rates her pain at a 8 out of 10 on a scale of 0 to 10. The physician orders morphine sulfate for pain.  

The order reads: MS 15mg IM, q1-3hrs PRN incisional pain. /Dr. Smith

You are preparing an injection of morphine. The order reads 15 mg IM. The stock morphine is 48mg/2mL. You will draw up ________mL (round to the nearest 10th)

  1. An order reads: Give 1000 mL over 8 hours at 20 gtt/mL. What is the flow rate in _______gtt/min? (Remember with IV's to round your answer to a whole number)

  1. What is the flow rate per hour of 1000mL to run over 6 hours? _____mL/hr.  

(Remember with IV's to round to the nearest whole number)

  1. The nurse has a patient with a respiratory infection and receives orders to start the antibiotic cephalexin (Keflex) 50 mg/kg/24 hours divided into two equal doses for an adolescent girl that weighs 88 lbs. The medication is provided in 500-mg tablets. How many tablets will the nurse provide with each dose?_____tablet

  1. 1:10 ::X:50 Determine the value of X

  1. The doctor orders Lactulose 30 mL po . How many tablespoons should you give?_____T

  1. Fill in the following prefix with its level of measurement.
      
    ____________tens (1/10)    

A) Deci-     B)Centi-   C) Milli- D)Kilo-

  1. You want heparin 300 units/kg body weight. You have heparin 20,000 units/mL. The patient weighs 185 pounds. How many mL of medication will you give. ______mL (Round to the nearest 10th.)
Answer #1

ANSWER.

1) Calculation of amount draw up.

       Formula = Desired dose/Dose in hand * Quantity.

    Here

         Desired dose (D) = 15 mg

         Dose in hand (H) = 48 mg

              Quantity (Q) = 2ml.

As per above formula

     Amount (ml) of drug draw up = D/H * Q

                                             = 15 / 48 * 2

                                             = 0.31 * 2

                                             = 0.6 ml.

2) Flow rate in gtt/min.

        Formula

           Flow rate (gtt/min) = Volume(ml) /Time(min) * Drop factor (gtt/ml).

     Here

          Volume(ml) = 1000 ml.

          Time(min) = 480 min (8hr).

           Drop factor = 20 gtt/ml.

    As per above formula

         Flow rate (gtt/min) = 1000 / 480 * 20

                                       = 2.08 * 20

                                       = 41.66

                                       = 42 gtt/min.

3) Calculation of ml/hr.

        Formula

             ml/hr = Volume(ml) / Time(hr).

    Here

        Volume(ml) = 1000 ml.

         Time(hr) = 6 hr.

As per above formula

           ml/hr = 1000 / 6

                    = 166.66

                    = 167 ml/hr.

4) Calculation of number of tablet required per dose.

           Ordered dose = 50 mg/kg/24 hrs (2 equal doses).

           Weight of the patient = 88 lbs.

     Convert lbs to kg for calculation.

              1 lbs = 0.454 kg.

      So

          Weight of the patient in kg = 88 * 0.454

                                                     = 39.95 kg.

       Dose required for a child weight 39.95 kg = 50 * 39.95

                                                                          = 1997.5 mg / 24 hr.

        Dose of available tablets = 500 mg/tablet.

     Number of tablets required per day = Dose needed / Available dose.

                                                                 = 1997.5 / 500

                                                                 = 3.99

                                                                 = 4 tablet / 24 hr.

      Number of tablet required per dose = 4/2 ( 2 equal doses per 24 hr).

                                                                = 2 tablets.

5) 1:10 = X:50 value of X.

            1/10 = X/50

            10X = 1* 50

           X = 1 * 50 / 10

                  X = 5

6) Calculation of tablespoon of lactulose should give.

         Ordered amount = 30 ml.

             1 tablespoon = 15 ml.

    So

        Tablespoon of lactulose required = 30 / 15

                                                              = 2 tablespoon.

7) A) Desi (0.1) tens = 1/10

8) Calculation of volume(ml) of drug required.

         Formula = Desired dose/Dose in hand * Quantity.

    Here

         Ordered dose = 300 unit/kg.

       Weight of the patient = 185 pounds.

      Convert pound to kg for calculation

                 1 pound = 0.454 kg.

         Weight of the patient in kg = 185 * 0.454

                                                    = 83.99 kg.

           Desired dose (D) = 300 * 83.99

                                       = 25197 unit.

           Dose in hand (H) = 20000 unit.

                Quantity (Q) = 1 ml.

As per above formula

        Volume(ml) of drug required = D/H * Q

                                                     = 25197 / 20000 * 1

                                                     = 1.3

                                                     = 1 ml.